Calculate the mass of Potassium ferricyanide required to prepare 10ml of 10mM potassium ferricyanide in aqueous 1.0M potassium nitrate solution
Answer #1

The mass of sodium ferricyanide is calculated:

g SF = M * V * MM = 0.01 M * 0.01 L * 422.39 g / mol = 0.0422 g

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.